# Floating Edge Bands in the Bernevig-Hughes-Zhang model with Altermagnetism

### v1

@@ I Introduction
Conventionally, edge bands connect (and merge into) the bulk conduction and valence bands, and exist only in a limited range of momentum space [1, 2, 3]. Recently, it has been found that edge bands can emerge over the entire Brillouin zone and may detach from the bulk bands, floating inside the band gap [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. This exotic band structure has been identified in various systems, including layered nonsymmorphic semimetals [4, 5, 6, 7, 8, 9] and obstructed atomic insulators [10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19], and attracted broad attention. With their unique properties, floating edge bands (FEB) are robust against back scattering and offer a platform for frequency multiplexing transport. They are also closely related to other intriguing topological phases, such as higher-order topological states [17, 15, 16, 13, 16, 14].
@@
Recently, a third type of collinear magnetic order called altermagnetism has been observed in various materials [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. This novel magnetic phase may be caused by Fermi liquid instabilities [48, 49, 50]. It can also emerge from crystal symmetries that alternate its spin polarization in both coordinate and momentum space while maintaining net zero magnetization [43, 44]. Unlike conventional ferromagnetism and antiferromagnetism, altermagnetism possesses non-relativistic spin splitting which anisotropically depends on momentum, e.g., analogous to $d$-wave pairing potential in superconductors [44]. These distinctive properties prompt several exotic phenomena [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], such as anomalous Hall effects [72, 73, 74, 75], giant tunneling magnetoresistance [76, 77], and finite-momentum Cooper pairing [53, 64, 78, 79, 80].
@@ II BHZ model with $d$-wave altermagnetism
where $M({\bf k})=M_{0}+2B(\cos k_{x}+\cos k_{y})$ with $M_{0}=M-4B$ and ${\bf k}=(k_{x},k_{y})$ being the Bloch wave vector in the plane. It is written in the basis $\Psi=(c_{\uparrow,a},c_{\uparrow,b},c_{\downarrow,a},c_{\downarrow,b})^{T}$ with subscripts $a$ and $b$ distinguishing two orbitals of opposite parity. $\bm{s}=(s_{1},s_{2},s_{3})$ are Pauli matrices for spins and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are Pauli matrices for the two orbitals. The first line in Eq. (1) is the original BHZ model, where the $A$ term accounts for orbital-momentum locking and the $M({\bf k})$ term is the mass term. The $J$ and $G$ terms describe the $d$-wave altermagnetic and Zeeman fields, respectively. Without loss of generality, we consider $A>0$  [^1] For a negative $A$, the topological invariants flip their signs.. We set the lattice constant to be unity. To illustrate the key results, we further assume that the two orbitals have opposite responses to the fields. However, we note that the main results extend beyond this unessential assumption, as we will show later [Sec. IV]. In the BHZ model, the two spin states are uncoupled. Thus, the Hamiltonian split into two $2\times 2$ blocks even in the presence of the altermagnetic and Zeeman fields in the $z$ direction. The two blocks can be related by
@@ III.1 Topological Numbers
Next, we show the emergence of FEB in the BHZ model under the altermagnetism. We first derive the topological invariants for the model. Since the model is decoupled into two blocks, we can treat them separately.
@@
$d_{3}=$ $M(\bm{k})+2J(\cos k_{y}-\cos k_{x})+G.$ (4)
The topological property of the two-band subblock Hamiltonian can be characterized by a Chern number defined on the first Brillouin zone [82]:
$C_{+}=-\frac{1}{4\pi}\int_{\text{1BZ}}d^{2}{\bf k}\frac{\bm{d}\cdot(\partial_{k_{x}}\bm{d}\times\partial_{k_{y}}\bm{d})}{d^{3}},$ (5)
where $d=\sqrt{\sum_{\nu=1}^{3}d_{\nu}^{2}}$. A straightforward calculation yields
@@
Fig. 1: (a,b) Topological phase diagrams classified by Chern number (blue/red for $C=\pm 1$, grey for $C=0$) and winding number $(W_{0},W_{\pi})$ for the (a) spin-up block and (b) spin-down block. Other parameters are $M=2$, $A=1$ and $B=1.4$. (c) Topological phase diagram of the full altermagnetic BHZ model as functions of altermagnetic $J$ and Zeeman field $G$ for $M=2$, $A=1$ and $B=1.4$. The floating edge band (FEB) phase is characterized by winding numbers $(W_{0},W_{\pi})=(-1,-1)$ for either spin. The suspending edge band (SEB) phase is characterized by the total winding numbers $(W_{0},W_{\pi})=(-1,-1)$ given by the summation of both spins. Solid (dashed) lines correspond to the phase boundaries of the spin-up (spin-down) block. The FEB phase contains regions I, II, and III corresponding to an FEB phase in one spin block while different topological phases in the other. (d) The same as (c) but for $B=-1$.
@@
Although the overall chiral symmetry is not preserved in the full Hamiltonian, at some specific $k_{x}$, i.e., $k_{x}=0$ and $k_{x}=\pi$, an emergent chiral symmetry $\mathcal{C}=\sigma_{1}$ appears. This is illustrated by $\{\mathcal{C},H_{0/\pi}\}=0$. This chiral symmetry allows us to classify the topological phases by winding number [83]. To obtain the winding number, we first perform a unitary transformation $U=e^{i\frac{\pi}{4}\sigma_{2}}$ to diagonalize the chiral symmetry operator. After the transformation, the chiral symmetry operator becomes $\tilde{\mathcal{C}}=U\mathcal{C}U^{-1}=\sigma_{3}$. Accordingly, the spin-up Hamiltonian transforms to
@@
Fig. 2: Spin-resolved density of states (DOS) along one boundary obtained by the surface Green’s function method [84]. (a) Spin-resolved DOS for $B=1.4$ and $J=G=0$. The system is in a quantum spin Hall phase with gapless helical edge states across the bulk gap. (b) Spin-resolved DOS for phase I in Fig. 1 ($B=1.4$, $J=1$, and $G=1$). (c) Spin-resolved DOS for phase II in Fig. 1 ($B=1.4$, $J=-2$, and $G=-1$). The spin-up edge bands overlap with the spin-down FEB around $k_{x}=0$. (d) Spin-resolved DOS for phase III in Fig. 1 ($B=1.4$, $J=-2$ and $G=-1$). The spin-up edge bands partially overlap with the spin-down FEB, while the spin-up bulk bands touch the FEB. (e) Spin-resolved DOS for $B=1.4$, $J=1$ and $G=4$. The spin-up FEB is completely detached from the bulk bands of both spins, forming a SEB phase. (f) Spin-resolved DOS for $B=-1$, $J=3$ and $G=1$. Edge bands of different spins are located in different $k_{x}$ regions. Other parameters for panels are $A=1$ and $M=2$.
@@ III.2 Diverse Topological Phases
By considering the combined topological phases of both spin blocks, we arrive at the phase diagram depicted in Fig. 1(c). The solid lines represent the phase boundaries of the spin-up block, while the dashed lines indicate those of the spin-down block. Remarkably, the FEB phase, characterized by winding numbers $(W_{0},W_{\pi})=(\pm 1,\pm 1)$ in either block, emerges even though the bulk system holds a zero Chern number. This phase extends across broad ranges of the parameters, as highlighted by the color-shaded areas in Fig. 1(c). When the spin-down block enters a topological regime, we can find an additional spin-down edge band around $k_{x}=0$. These bands can either form a quantum spin Hall regime [see Fig. 2(b)] or overlap with the spin-up FEB [see Fig. 2(c)]. These two phases are labelled as phases I and II in Fig. 1(c), respectively. Moreover, the spin-down block can transit into another topological phase, leading to another phase marked as phase III in Fig. 1(c). In this phase, the spin-down edge bands may overlap with the FEB from the spin-up block and its bulk bands may touch the FEB [see Fig. 2(d)]. Thus, to obtain edge bands completely detached from the bulk bands, one block, say the spin-down block, must be driven into a trivial phase. These topological phases, where the bulk and edge bands are completely separated for the full system [see Fig. 2(e)], are denoted as SEB in Fig. 1(c).
@@
The appearance of the FEB phase by altermagnetism does not depend on a specific range of $M$ and $B$. We can also reach the FEB phase from the trivial topological phase at zero fields. For example, in the case with $M_{0}>-4B>0$, the SEB phase from the spin-down block can be obtained by introducing a Zeeman field $G$ to invert the bands at $k_{x}=0$ and then adding the altermagnetic field $J$ to invert the bands at $k_{x}=\pi$, as shown in Fig. 1(d). Note that in this case, a constant Zeeman field is also crucial for the formation of the FEB. Another interesting topological phase shown in Fig. 2(f) appears when total winding number $(W_{0},W_{\pi})=(\pm 1,\pm 1)$ is assembled by $(W_{0}^{+},W_{\pi}^{+})=(\pm 1,0)$ and $(W_{0}^{-},W_{\pi}^{-})=(0,\pm 1)$. Though such a phase shares the same total winding number with the FEB phase, its edge bands do not extend to the whole Brillouin zone. The spin-up and spin-down edge bands may patch together to form a fragmented "FEB". It is worth noting that when the edge bands of different spins do not overlap in $k_{x}$ space, the edge states on the same edge will not interact with each other even in the presence of superconductivity and in-plane magnetization [Sec. IV].

### v3

@@ I Introduction
Conventionally, edge bands bridge (and merge into) the bulk conduction and valence bands, and exist only in a limited range of momentum space [1, 2, 3]. Recently, it has been found that edge bands can emerge over the entire Brillouin zone and may detach from the bulk bands, floating inside the band gap [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. This exotic band structure has been identified in various systems, including layered nonsymmorphic semimetals [4, 5, 6, 7, 8, 9] and obstructed atomic insulators [10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19], and attracted broad attention. With their unique properties, floating edge bands (FEB) are robust against back scattering and offer a platform for frequency multiplexing transport [20]. They are also closely related to other intriguing topological phases, such as higher-order topological states [21, 17, 22, 15, 16, 13, 16, 14].
@@
Recently, a third type of collinear magnetic order called altermagnetism has been observed in various materials [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. This novel magnetic phase may be caused by Fermi liquid instabilities [53, 54, 55]. It can also emerge from crystal symmetries that alternate its spin polarization in both coordinate and momentum space while maintaining net zero magnetization [47, 48]. Unlike conventional ferromagnetism and antiferromagnetism, altermagnetism possesses non-relativistic spin splitting which anisotropically depends on momentum, e.g., analogous to $d$-wave pairing potential in superconductors [48]. These distinctive properties prompt several exotic phenomena [21, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76], such as anomalous Hall effects [77, 78, 79, 80], giant tunneling magnetoresistance [81, 82], and zero-field finite-momentum Cooper pairing [58, 69, 83, 84, 85].
@@ II BHZ model with $d$-wave altermagnetism
where $M({\bf k})=M_{0}+2B(\cos k_{x}+\cos k_{y})$ with $M_{0}=M-4B$ and ${\bf k}=(k_{x},k_{y})$ being the Bloch wave vector in the plane. It is written in the basis $\Psi=(c_{\uparrow,a},c_{\uparrow,b},c_{\downarrow,a},c_{\downarrow,b})^{T}$ with subscripts $a$ and $b$ distinguishing two orbitals of opposite parity. $\bm{s}=(s_{1},s_{2},s_{3})$ are Pauli matrices for spins and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are Pauli matrices for the two orbitals. The first line in Eq. (1) is the original BHZ model, where the $A$ term accounts for orbital-momentum locking and the $M({\bf k})$ term is the mass term. The $J$ and $G$ terms describe the $d$-wave altermagnetic and Zeeman fields, respectively. While the Zeeman field $G$ can be introduced by applying external magnetic fields, the altermagnetic field can be achieved via proximity effect by coupling the BHZ system to an altermagnetic material. Alternatively, the altermagnetism can be realized and controlled using antiferroelectricity [86]. Without loss of generality, we consider $A>0$  [^1] For a negative $A$, the topological invariants flip their signs.. We set the lattice constant to be unity. To illustrate the key results, we further assume that the two orbitals have opposite responses to the fields. However, we note that the main results extend beyond this unessential assumption, as we will show later [Sec. IV]. In the BHZ model, the two spin states are uncoupled. Thus, the Hamiltonian split into two $2\times 2$ blocks even in the presence of the altermagnetic and Zeeman fields in the $z$ direction. The two blocks can be related by
@@ III.1 Topological Numbers
Next, we show the emergence of FEB in the BHZ model under the altermagnetic field. We first derive the topological invariants for the model. Since the model is decoupled into two blocks, we can treat them separately.
@@
$d_{3}=$ $M(\bm{\mathrm{k}})+2J(\cos k_{y}-\cos k_{x})+G.$ (4)
In the model, there is no TRS, particle-hole symmetry, or chiral symmetry, thus it belongs to class A [89] in the $\mathbb{Z}$ classification. The topological property of the two-band subblock Hamiltonian can be characterized by a Chern number defined on the first Brillouin zone [88]:
$C_{+}=-\frac{1}{4\pi}\int_{\text{1BZ}}d^{2}{\bf k}\frac{\bm{d}\cdot(\partial_{k_{x}}\bm{d}\times\partial_{k_{y}}\bm{d})}{|\bm{d}|^{3}},$ (5)
where $|\bm{d}|=\sqrt{\sum_{\nu=1}^{3}d_{\nu}^{2}}$. A straightforward calculation yields
@@
Fig. 1: (a) Topological phase diagram in the spin-up block classified by winding number $(W_{0}^{+},W_{\pi}^{+})$ and Chern number $C_{+}$ (shaded in blue/red for $C_{+}=\pm 1$, and grey for $C_{+}=0$). (b) The same as (a) but for the spin-down block. Other parameters for both (a) and (b) are $M=2$, $A=1$, and $B=1.4$. (c) Topological phase diagram of the full altermagnetic BHZ model as functions of altermagnetic $J$ and Zeeman field $G$, with parameters $M=2$, $A=1$, and $B=1.4$. The FEB phase is characterized by winding numbers $(W_{0}^{\pm},W_{\pi}^{\pm})=(-1,-1)$ for either spin ("+" or "-"). The SEB phase is characterized by the total winding numbers $(W_{0},W_{\pi})=(-1,-1)$, which is the sum over both spins. Solid (dashed) lines correspond to the phase transition line of the spin-up (-down) block. The FEB phase contains regions I, II, and III, corresponding to a FEB phase in one spin block while different topological phases (shown in Figs. 2(b-d)) in the other spin block. (d) The same as (c) but for $B=-1$. QSH and TP I denote different topological phases shown in Figs. 2(a) and (f), respectively.
@@
[…]
@@ III.2 Diverse Topological Phases
Fig. 2: Spin-resolved density of states (DOS) along one boundary obtained by the surface Green’s function method [93]. (a) Spin-resolved DOS for $B=1.4$ and $J=G=0$. The system is in a quantum spin Hall (QSH) phase with gapless helical edge states across the bulk gap. (b) Spin-resolved DOS for phase I in Fig. 1 ($B=1.4$, $J=1$, and $G=1$). (c) Spin-resolved DOS for phase II in Fig. 1 ($B=1.4$, $J=4$, and $G=6$). The spin-up edge bands overlap with the spin-down FEB around $k_{x}=0$. (d) Spin-resolved DOS for phase III in Fig. 1 ($B=1.4$, $J=2$ and $G=1$). The spin-up edge bands partially overlap with the spin-down FEB, while the spin-up bulk bands touch the FEB. (e) Spin-resolved DOS for $B=1.4$, $J=1$ and $G=4$. The spin-up FEB is completely detached from the bulk bands of both spins, forming a SEB phase. (f) Spin-resolved DOS for $B=-1$, $J=3$ and $G=1$. Edge bands of different spins are located in different $k_{x}$ regions. This topological phase is denoted as TP I in Fig. 1 (d). Other parameters for panels are $A=1$ and $M=2$.
@@
By considering the combined topological phases of both spin blocks, we arrive at the phase diagram depicted in Fig. 1(c). The solid lines represent the phase boundaries of the spin-up block, while the dashed lines indicate those of the spin-down block. Remarkably, the FEB phase, characterized by winding numbers $(W_{0}^{\pm},W_{\pi}^{\pm})=(\pm 1,\pm 1)$ in either block ("+" or "-"), emerges even though the bulk system holds a zero Chern number. This phase extends across broad ranges of the parameters, as highlighted by the color-shaded areas in Fig. 1(c). When one of the spins (e.g. spin-up) enters the FEB phase while the other (spin-down) enters a topological regime with low-energy modes at $k_{x}$=0, the spin-down band can either form a quantum spin Hall regime with the spin-up FEB [see Fig. 2(b)] or overlap with the spin-up FEB [see Fig. 2(c)]. These two phases are labelled as phases I [characterized by ($W_{0},W_{\pi}$)=(±$\pm 1,\pm 1$) for one spin and ($W_{0},W_{\pi}$)=(±$\pm 1,0$) for the other] and phase II [characterized by ($W_{0},W_{\pi}$)=(±$\pm 1,\pm 1$) for one spin and ($W_{0},W_{\pi}$)=($\mp 1,0$) for the other] in Fig. 1(c), respectively. The overlapping edge states can be split, for instance, by a spectrum asymmetry term $2\delta J(\cos k_{y}-\cos k_{x})s_{3}\sigma_{0}$. Moreover, the spin-down block can transit into another topological phase, leading to another phase marked as phase III [characterized by ($W_{0},W_{\pi}$)=(±$\pm 1,\pm 1$) for one spin and ($W_{0},W_{\pi}$)=(0,0) for the other] in Fig. 1(c). In this phase, the spin-down edge bands may overlap with the FEB from the spin-up block and its bulk bands may touch the FEB [see Fig. 2(d)]. Thus, to obtain edge bands completely detached from the bulk bands, one block, say the spin-down block, must be driven into a trivial phase. These topological phases, where the bulk and edge bands are completely separated for the full system [see Fig. 2(e)], are denoted as SEB in Fig. 1(c).
@@
The appearance of the FEB phase by altermagnetism does not depend on a specific range of $M$ and $B$. We can also reach the FEB phase from the trivial topological phase at zero fields. For example, in the case with $M_{0}>-4B>0$, the SEB phase from the spin-down block can be obtained by introducing a Zeeman field $G$ to invert the bands at $k_{x}=0$ and then adding the altermagnetic field $J$ to invert the bands at $k_{x}=\pi$, as shown in Fig. 1(d). Note that in this case, a constant Zeeman field is also crucial for the formation of the FEB. As the altermagnetic term $2J(\cos k_{y}-\cos k_{x})$ takes different signs at $k_{x}=0$ and $k_{x}=\pi$, it alone cannot induce band inversions at both $k_{x}=0$ and $k_{x}=\pi$ to form the FEB phase. Another interesting topological phase shown in Fig. 2(f) appears when total winding number $(W_{0},W_{\pi})=(\pm 1,\pm 1)$ is assembled by $(W_{0}^{+},W_{\pi}^{+})=(\pm 1,0)$ and $(W_{0}^{-},W_{\pi}^{-})=(0,\pm 1)$. Though such a phase shares the same total winding number with the FEB phase, its edge bands do not extend to the whole Brillouin zone. The spin-up and spin-down edge bands may patch together to form a fragmented "FEB" with non-zero Chern number $C=C_{+}+C_{-}=-2$. It is worth noting that when the edge bands of different spins do not overlap in $k_{x}$ space, the edge states on the same edge will not interact with each other even in the presence of superconductivity and in-plane magnetization [Sec. IV]. In the appendix, we present the explicit evolution of these states by changing the parameters $G$ and $J$ in the phase diagram.
@@ Acknowledgements.
Appendix A Evolution of Topological Phases
We present the evolution of the topological phases in detail. Transitions between distinct topological phases are marked by gap closings at either $k_{x}=0$ or $k_{x}=\pi$. To illustrate this, we examine the cases of $-4B<M_{0}<0$ and $M_{0}>-4B>0$, respectively, where we define $M_{0}=M-4B$ for brevity.
Fig. 5: (a) The evolution path $A\to B\to C\to D\to E\to F\to A$ that traces the transitions between topological phases in the phase diagram. (b,c) The band structure evolution along the loop path marked in (a). Phase transitions occur when the bulk gap closes at (b) $k_{x}=0$ or (c) $k_{x}=\pi$. Other parameters are $M=2$, $B=1.4$, $A=1$.
Fig. 6: Spin-resolved DOS for (A) QSH phase with $(J,G)=(0,0)$, (B) QAH phase with $(J,G)=(0,3)$, (C) SEB phase with $(J,G)=(1,3)$, (D) FEB II phase with $(J,G)=(3,3)$, (E) FEB III phase with $(J,G)=(3,1)$, and (F) FEB I phase with $(J,G)=(1,1)$. The other panels display the gap-closing points at transitions between these phases. Other parameters are $M=2$, $B=1.4$, and $A=1$.
The case of $-4B<M_{0}<0$: The phase evolution along the loop path $A\to B\to C\to D\to E\to F\to A$, marked in Fig. 5(a), is illustrated in Figs. 5(b) and 5(c). The phase transitions occur when the bulk gap closes at $k_{x}=0$ [see Fig. 5(b)] or $k_{x}=\pi$ [see Fig. 5(c)]. To be specific, we start with the quantum spin Hall (QSH) phase (point A) at $J=G=0$ in Fig. 6A and introduce the Zeeman field $G$. The bulk gap closes when the path crosses the phase boundary at $G=M$, removing the edge state for one spin and entering the quantum anomalous Hall (QAH) phase (point B). Next, we turn on the altermagnetism $J$, which closes the gap at the phase boundary $G+4J=-M_{0}$, and enter the SEB phase (point C), creating an additional low-energy spin-up edge state at $k_{x}=\pi$. From the SEB phase, as we further increase the altermagnetic strength, the gap closes when the path crosses the phase boundary at $G-4J=M_{0}$, entering the FEB II phase (point D) with spin-down edge modes overlapping with the spin-up FEB. Subsequently, we reduce the strength of the Zeeman field. The bulk gap closes when the path crosses the phase boundary at $G=M$, removing the spin-down edge modes around $k_{x}=0$ and transitioning the system into the FEB III phase (point E). Reducing the altermagnetic field further results in a gap closing a $G-4J=M_{0}$, creating spin-down edge modes at $k_{x}=0$ and entering the FEB I phase (point F). Finally from point F in the FEB I phase, we turn off both altermagnetic and Zeeman fields. The bulk gap closes when the path crosses the phase boundary at $G+4J=-M_{0}$, removing the spin-up edge modes at $k_{x}=\pi$ and returning to the QSH phase (point A).
Fig. 7: (a) The evolution path $A\to B\to C\to D\to E\to A$ that traces the transitions between topological phases in the phase diagram. (b,c) The band structure evolution along the loop path marked in (a). Phase transitions occur when the bulk gap closes at (b) $k_{x}=0$ or (c) $k_{x}=\pi$. Other parameters are $M=2$, $B=-1$, $A=1$.
Fig. 8: Spin-resolved DOS for (A) trivial phase with $(J,G)=(0,0)$, (B) TP II phase with $(J,G)=(0,4)$, (C) SEB phase with $(J,G)=(1.5,4)$, (D) FEB II phase with $(J,G)=(4,4)$, and (E) TP I phase with $(J,G)=(3,0)$. The other panels display the gap-closing points at transitions between these phases. Other parameters are $M=2$, $B=-1$, and $A=1$.
The case of $M_{0}>-4B>0$: The phase evolution along the loop path $A\to B\to C\to D\to E\to A$, marked in Fig. 7(a), is illustrated in Figs. 7(b) and 7(c) The phase transitions occur when the bulk gap closes at $k_{x}=0$ [see Fig. 7(b)] or $k_{x}=\pi$ [see Fig. 7(c)]. To be specific, we start with the trivial phase (point A) at $J=G=0$ in Fig. 8A and introduce the Zeeman field $G$. The bulk gap closes when the path crosses the phase boundary at $G=M$, creating spin-down edge modes at $k_{x}=0$ and entering the topological (TP) II phase (point B). Next, we turn on the altermagnetism to close the gap when the path crosses the phase boundary at $G+4J=M_{0}$, further creating another spin-down edge mode at $k_{x}=\pi$ and transitioning the system into the SEB phase (point C). From the SEB phase, we then increase the altermagnetic strength, the gap closes when the path crosses the phase boundary at $G-4J=-M_{0}$, transitioning the system into the FEB II phase (point D) by creating spin-up edge modes at $k_{x}=0$. Subsequently, we reduce the strengths of both magnetic fields. The gap closes when the path crosses the phase boundary at $G=M$, removing spin-down edge modes around $k_{x}=0$ and transitioning the system into the TP I phase (point E). Finally, we turn off the altermagnetic field and close the gap when the path crosses the phase boundary at $4J=M_{0}$, removing the edge modes and returning the system to the trivial phase (point A).